OpenStudy (anonymous):

You kick a soccer ball with a speed of 12 m/s at an angle of 21. How long does it take the ball to reach the top of its trajectory?

OpenStudy (samgrace):

using a diagram resolve for x and y directions, x = 12cos21 y = 12sin21 using v = u + at, solve for t with v = 0 and u = 12sin21 to give time it takes to reach the height of... v^22 = u^2 + 2as, substitute y = 12sin21 = u, and set v =0, a = -9.81 m/s and solve for s, this will give you it's maximum height.

OpenStudy (anonymous):

you can determine the starting vertical component of velocity from vy0 = 12 sin 21 and note that gravity reduces the velocity by -9.8m/s^2 so it stops when 0 = vy0 - gt = 12 sin 21 - (9.8 m/s) t and solve for t